# Quantitative Comparison of Fine-Tuning Techniques for Pretrained Latent Diffusion Models in the Generation of Unseen SAR Image Concepts

### v1

@@ 3 Dataset Creation
Pre-processing Raw Data
Training Dataset Creation
Our final dataset consists of refocused and resampled SAR images stored as complex-valued matrices. For training purposes, we work on amplitude images, whose pixel values follow a Rayleigh distribution — unbounded and highly skewed, with a small proportion (1–3%) of very high-intensity scatterers. These bright pixels are critical as they correspond to strong reflectors such as buildings or metallic structures. For visualization and learning stability, we apply the following normalization:
$A(r,y)_{\text{norm}}=\frac{A(r,y)}{\mu+3\cdot\sigma}$ (1)
where $A(r,y)$ is the amplitude value at coordinates $(r,y)$, and $\mu$ and $\sigma$ are the mean and standard deviation of the amplitude image, respectively. Then, we clip values between 0 and 1, ensuring that 98% of the values fall within this interval. This results in all pixels beyond the threshold are "flattened" at exactly 1. This creates an artificial saturation at pixel value 1.
The normalized images are then cropped into standardized patches of size 1024 × 1024 pixels. For a subset of these patches, we created geo-aligned SAR–optical image pairs using optical imagery from the IGN database. Textual descriptions were automatically generated for the optical images using the foundation model CogVLM2 [26].
@@ 4.1 Stable Diffusion framework
The model operates in a latent space rather than directly in pixel space. Text–image pairs are processed independently into compact latent representations. First, a Variational Autoencoder (VAE) encodes an image $x$ into a latent representation $z$:
@@
As shown in the Figure 3, the model learns by noising training data (i.e. VAE Encoder output), through the successive addition of Gaussian noise (forward process). Then the model is used reverse this process (reverse process) to turn noise back into data by removing the noise added during each diffusion step. More specifically, the latent $z_{t}$ is the output of the VAE encoder from the input image $x_{t}$, and noise is added to simulate a step in the forward diffusion process, governed by a scheduler that defines the noise distribution at each $t$. More specifically, the forward diffusion process, which adds noise to an input data $z_{0}$ over $T$ timesteps by adding Gaussian noise, is defined as:
@@
During training (Figure 4), the model receives a corrupted latent $z_{t}$ for a randomly sampled timestep $t\in[0,1000]$ and is tasked with predicting the corresponding $\epsilon_{t}$. The loss is computed over multiple such noise levels per batch and epoch.
@@
During inference (generation), the process begins from a pure Gaussian noise sample $z_{T}$. The UNet iteratively denoises this latent through $T$ steps, each conditioned on the text prompt. Although the model was trained on a schedule of $T=1000$ steps, it is common in practice to use only 25 to 50 steps for generation. At each step $t$, the model predicts $\epsilon_{t}$, and uses the reverse schedule to approximate $z_{t-1}$. After the final step, the latent $z_{0}$ is decoded by the VAE to produce the final image $x_{0}$.
@@ 4.2 Training approaches and parameterization
    LoRA-based tuning of the text encoders may be sufficient, as the semantic embedding space for textual prompts remains relatively stable and may not require extensive retraining across modalities.
In Section 5, we empirically evaluate these strategies across several model configurations. We use both semantic and statistical metrics to determine the most effective fine-tuning paradigm for generating realistic and coherent SAR images from textual descriptions.
@@ 4.3 Evaluation of generated SAR images
To assess the realism of the generated SAR images, we use a test dataset composed of triplets: [captions, labels, real SAR images]. Since SAR intensity distributions vary significantly depending on the type of scene, we perform a label-specific evaluation to account for this variability. The labels correspond to semantic scene categories — forest, field, city, airport, seacoast, port, mountains, beach, industrial, and residential — derived from a manually created keyword dictionary. For each label, the associated test captions are used as prompts to generate synthetic SAR images. In total, we generate 30 images per label across 11 categories, resulting in 330 generated images used for evaluation and comparison across model configurations.
@@
To compare the amplitude distributions between real and generated SAR images, we flatten each image into a one-dimensional array of pixel amplitudes. As described in Section 3, a normalization step is applied during pre-processing, and pixel values are clipped to the range [0, 1], introducing an artificial saturation effect at the upper bound.
To enable accurate statistical comparisons using the Kullback–Leibler (KL) divergence, we first exclude all pixel values strictly above 0.999, corresponding to saturated pixels. We then compute the proportion of saturated pixels separately and renormalize the histograms over the remaining values so that the probability density integrates to 1.
@@
Beyond statistical similarity, we also evaluate how well the generated SAR images align semantically with their conditioning prompts. To this end, we fine-tuned a CLIP ViT-L/14 model, on a separate dataset containing SAR image–caption pairs, using a batch size of 100, with the goal of embedding both modalities into a common latent space adapted to radar imagery.
To quantify alignment, we adopt two complementary evaluation strategies. First, we compute the ranking score, which measures how well each image is matched to its correct caption among a set of other texts. For each batch of $N=16$ image–text pairs, we extract normalized image embeddings $f_{\text{img}}(x_{i})$ and text embeddings $f_{\text{text}}(t_{j})$, and compute the cosine similarity matrix $S\in\mathbb{R}^{N\times N}$:
@@
We apply a softmax normalization across each row of the matrix to interpret the values as match probabilities. For each image $x_{i}$, we compute the rank $r_{i}$ of its corresponding ground-truth caption $t_{i}$ within the list of possible captions. We report the mean rank $r_{\mu}$, median rank and rank variance $r_{\sigma}$ over the evaluation set:
@@
In addition to ranking-based evaluation, we compute the cosine similarity between each generated image and its prompt, using the same fine-tuned SAR CLIP model. For each of the 11 semantic labels, we average these similarity scores across all generated samples to obtain a per-label, per-model alignment metric. The results are presented as heatmaps, where higher values indicate stronger text–image coherence. As a reference baseline, we compute the same similarity scores between real SAR images and their corresponding captions using the same CLIP model. These values are included at the bottom of each heatmap for visual comparison.
This two-pronged evaluation approach allows us to assess both relative ranking performance (i.e., how uniquely matched each prompt is to its image) and absolute similarity (i.e., how close the embedding vectors are).
@@
To evaluate the structural realism of the SAR images generated by our model, we compute textural indicators derived from the Gray-Level Co-occurrence Matrix (GLCM), following the classical method proposed by Haralick et al. [27].
We base this analysis on the same set of 330 labeled real and generated SAR images. From each image, we extract homogeneous patches of size 64 × 64 pixels, using segmentation masks inferred by the Segment Anything Model (SAM) [28]. Specifically, we identify the largest mask in each image and apply a sliding kernel to extract patches that are fully contained within that region. For each patch, we compute the GLCM.
@@
    Entropy: $-\sum_{l,k}\text{GLCM}(l,k)\log(\text{GLCM}(l,k)+\epsilon)$
These features characterize various aspects of image texture such as spatial regularity, smoothness, contrast, and randomness. We compute them across multiple orientations $\theta$ and distances $d$ to assess rotation-invariant properties.
@@ 5.1 Importance of the noise offset
When fine-tuning the model on SAR images, we introduce a small noise offset during the forward diffusion process. At each training step, the standard Gaussian noise $\varepsilon\sim\mathcal{N}(0,1)$ is perturbed by an additional random term, defined as:
@@
This modification effectively shifts the noise distribution to $\mathcal{N}(0,1+\gamma^{2})$, introducing a variability per sample and per channel without altering the spatial structure of the noise. For SAR imagery, which exhibits heavy-tailed amplitude distributions and pronounced contrast between bright scatterers (e.g., buildings, ships) and low-return regions (e.g., calm water), this additional stochasticity helps the model better capture the full dynamic range.
@@
Table 2: Comparison of trainings - at epoch 8 - with and without noise offset and LoRA (r: rank, a: alpha).
@@
Figure 6: Comparison of image generated during training - at epoch 8 - ith the same prompt: "A satellite view of a port with a boat in the water and a forest nearby." (see more examples in Appendix  A.2)
As illustrated in Figure 6, the absence of a noise offset leads to lower-contrast images with dull gray tones, failing to reproduce key SAR-specific features—such as the strong reflectivity contrast between land and sea. In contrast, introducing a small noise offset improves dynamic range modeling and enhances the physical realism of textures, as reflected by both visual inspection and the significant drop in KL divergence (Table 2). Based on these results, we apply a noise offset by default in all subsequent training runs.
@@ 5.2 Study on the UNet, TE1 and TE2
In this experiment, we compare different fine-tuning strategies for each major module of the SDXL architecture, namely, the UNet backbone, Text Encoder 1 (TE1), and Text Encoder 2 (TE2), to assess their relative importance and degree of independence. Each module is either fully fine-tuned (i.e., all weights are updated) or fine-tuned using Low-Rank Adaptation (LoRA) adapters.
[…]
@@
Refer to caption Refer to caption Real SAR Image image real zoom Refer to caption Refer to caption soleil-up-7 Image zoom Refer to caption Refer to caption smile-road-5 Image zoom
Figure 7: Comparison of image real and generated - at epoch 8 - with the same prompt: "A satellite view of a dynamic city with several buildings and a network of roads."
As shown in Table 3, full fine-tuning of the UNet consistently improves alignment and realism, as reflected in lower CLIP rank mean and variance. Although the model smile-road-5 achieves the best KL divergence score—indicating good statistical similarity to real SAR distributions—its high rank variance and weak alignment suggest instability. Indeed, visual inspection (Figure 7) reveals that this model sometimes generates unrealistic features, such as distorted textures or handwritten-like artifacts, which do not match the prompt.
In contrast, the model soleil-up-7 delivers both low CLIP rank and strong visual coherence. This configuration uses full UNet fine-tuning combined with LoRA on both text encoders, preserving prompt alignment while allowing the network to learn SAR-specific spatial patterns. Additional examples are shown in Appendix A.2, Figure 18.
To better understand which parts of the UNet contribute most to learning, we analyze the magnitude of parameter updates relative to the pretrained weights. As shown in Appendix A.1 (Figure 17), the largest changes are observed in the first ResNet blocks of the downsampling path and the final layers of the upsampling path. This suggests that early convolutional layers are critical for encoding SAR-specific noise and structure, further justifying the need for full fine-tuning of the UNet.
5.3 Study on LoRA Rank and Alpha
In this section, we investigate the impact of varying LoRA rank and scaling factor (alpha) on the performance of the fine-tuned model. All configurations use full fine-tuning of the UNet (F), while LoRA is applied to both text encoders (TE1 and TE2) with different rank–alpha pairs. Results are summarized in Table 4.

### v2

@@ 3 Dataset Creation
Pre-processing Raw Data
Training Dataset Creation
Our final dataset consists of refocused and resampled SAR images stored as complex-valued matrices. For training purposes, we work on amplitude images, whose pixel values follow a Rayleigh distribution — unbounded and highly skewed, with a small proportion (1–3%) of very high-intensity scatterers. These bright pixels are critical as they correspond to strong reflectors such as buildings or metallic structures. For visualization and learning stability, we apply the following normalization, which is a common practice in visual interpretation of SAR images to stabilize their dynamic range and improve the quality of visualization:
$A(r,y)_{\text{norm}}=\frac{A(r,y)}{\mu+3\cdot\sigma}$ (1)
where $A(r,y)$ is the amplitude value at coordinates $(r,y)$, and $\mu$ and $\sigma$ are the mean and standard deviation of the amplitude image, respectively. Values are clipped between 0 and 1 to have 98% of the values that fall in this interval. Thus, all pixels beyond the threshold are "flattened" at exactly 1. This creates an artificial saturation at pixel value 1.
However, it is important to note that this approach is a compromise. If the threshold is set too high to preserve the information of the strongest scatterers, it could cause all the weaker scatterers to be compressed into the same range of values, effectively losing the distinction between them. This would result in a loss of detail for the weaker scatterers, which are often important for accurate interpretation. Therefore, it’s essential to find a balance between preserving the information of the strongest reflectors (like buildings or metallic structures) while also maintaining enough precision to differentiate the weaker scatterers, which can be critical in some applications.
The normalized images are then cropped into standardized patches of size 1024 × 1024 pixels. For a subset of these patches, we created geo-aligned SAR–optical image pairs using optical imagery from the IGN database. Textual descriptions were automatically generated for the optical images using the foundation model CogVLM2 Hong et al. (2024).
@@ 4.1 Stable Diffusion framework
Stable Diffusion XL model operates in a latent space rather than directly in pixel space. Text–image pairs are processed independently into compact latent representations. First, a Variational Autoencoder (VAE) encodes an image $x$ into a latent representation $z$:
@@
As shown in the Figure 3, the model learns by noising training data (i.e. VAE Encoder output), through the successive addition of Gaussian noise (forward process). Then the model reverses this process (reverse process) to turn noise back into data by removing the noise added during each diffusion step. More specifically, the latent $z_{t}$ is the output of the VAE encoder from the input image $x_{t}$, and noise is added to simulate a step in the forward diffusion process, governed by a scheduler that defines the noise distribution at each $t$. More specifically, the forward diffusion process, which adds noise to an input data $z_{0}$ over $T$ timesteps by adding Gaussian noise, is defined as:
@@
During training (Figure 4), the model receives a corrupted latent $z_{t}$ for a randomly sampled timestep $t\in[0,1000]$ and is tasked with predicting the corresponding $\epsilon_{t}$. The loss is computed over multiple such noise levels per batch and epoch. Thus, to improve model performance without disrupting the pretrained knowledge, focusing on the final stages of the reverse diffusion process—corresponding to the earlier timesteps of the forward process during training—may offer a more effective refinement of the SAR image generation.
@@
By randomly choosing timesteps over a batch during training, the model is exposed to a diverse set of degradation levels, which enables it to better capture the underlying patterns across different stages of the reverse diffusion process. Therefore, it is crucial to fine-tune the model over several epochs, ensuring that the model sees all the data multiple times. This also guarantees that different categories of data, such as fields, forests, cities, or seacoasts, are seen at various timesteps within the range $t\in[0,1000]$ during each training epoch.
During inference (generation), the process begins from a pure Gaussian noise sample $z_{T}$. The UNet iteratively denoises this latent through $T$ steps, each conditioned on the text prompt. Although the model was trained on a schedule of $T=1000$ steps, it is common in practice to use only 25 to 50 steps for generation. At each step $t$, the model predicts $\epsilon_{t}$ and uses the reverse schedule to approximate $z_{t-1}$. After the final step, the latent $z_{0}$ is decoded by the VAE Decoder to produce the final image $x_{0}$.
@@ 4.2 Training approaches and parameterization
    LoRA-based tuning of the text encoders may help to preserve the model’s base language semantic knowledge, including spatial arrangements and object relationships, while adapting it to SAR imagery.
In Section 5, we empirically evaluate these strategies across several model configurations. We use both semantic and statistical metrics to determine the most effective fine-tuning techniques for generating realistic and coherent SAR images from textual descriptions.
@@ 4.3 Evaluation of generated SAR images
To assess the realism of the generated SAR images, we use a test dataset composed of triplets: [captions, labels, real SAR images]. Since SAR amplitude pixel values distributions vary significantly depending on the type of scene, we perform a label-specific evaluation to account for this variability. The labels correspond to semantic scene categories — forest, field, city, airport, seacoast, port, mountains, beach, industrial, and residential — derived from a manually created keyword dictionary. For each label, the associated test captions are used as prompts to generate synthetic SAR images. In total, we generate 30 images per label across 11 categories, resulting in 330 generated images used for evaluation and comparison across model configurations.
@@
To compare the amplitude distributions between real and generated SAR images, we flatten each image into a one-dimensional array of pixel amplitudes. As described in Section 3, a normalization factor is applied during pre-processing, and pixel values are clipped to the range [0, 1], introducing an artificial saturation peak at the upper bound.
To enable accurate statistical comparisons using the Kullback–Leibler (KL) divergence, we first exclude all pixel values corresponding to the saturated pixels (corresponding to 3% maximum). We then compute the proportion of saturated pixels separately and renormalize the histograms over the remaining values so that the probability density integrates to 1.
@@
Beyond statistical similarity, we also evaluate how well the generated SAR images align semantically with their conditioning prompts. To this end, we fine-tuned a CLIP ViT-L/14 model on a separate dataset containing SAR image–caption pairs, using a batch size of 100, with the goal of embedding both modalities into a common latent space adapted to radar imagery.
To quantify alignment, we adopt two other evaluation strategies. First, we compute the ranking score, which measures how well each image is matched to its correct caption among a set of other texts. For each batch of $N=16$ image–text pairs, we extract normalized image embeddings $f_{\text{img}}(x_{i})$ and text embeddings $f_{\text{text}}(t_{j})$, and compute the cosine similarity matrix $S\in\mathbb{R}^{N\times N}$:
@@
We apply a softmax normalization across each row of the matrix to interpret the values as match probabilities. For each image $x_{i}$, we compute the rank $r_{i}$ of its corresponding ground-truth caption $t_{i}$ within the list of possible captions. We report the mean rank $r_{\mu}$, median, and variance ranks $r_{\sigma}$ over the evaluation set:
@@
In addition to ranking-based evaluation, we compute the cosine similarity between each generated image and its prompt, using the same fine-tuned SAR-CLIP model. For each of the 11 semantic labels, we average these similarity scores across all generated samples to obtain a per-label, per-model alignment metric. The results are presented as heatmaps, where higher values indicate stronger text–image coherence. As a reference baseline, we compute the same similarity scores between real SAR images and their corresponding captions using the same CLIP model. These values are included at the bottom of each heatmap for visual comparison.
These evaluation approaches allows us to assess both relative ranking performance (i.e., how uniquely matched each prompt is to its image) and absolute similarity (i.e., how close the embedding vectors are).
@@
To evaluate the textural realism of the SAR images generated by our model, we compute textural indicators derived from the Gray-Level Co-occurrence Matrix (GLCM), following the classical method proposed by Haralick et al. Haralick et al. (1973). This method is well-suited for SAR texture analysis, as it allows us to assess directional patterns and spatial relationships in the generated images.
We base this analysis on the same set of 330 labeled real and generated SAR images. From each image, we extract homogeneous patches of size 64 × 64 pixels, using segmentation masks inferred by the Segment Anything Model (SAM) Kirillov et al. (2023). Specifically, we identify the largest mask in each image and apply a sliding kernel to extract patches that are fully contained within that region. For each patch, we compute the GLCM.
@@
    Entropy: $\sum_{l,k}\text{GLCM}(l,k)\log(\text{GLCM}(l,k)+\epsilon)$
These features characterize various aspects of image texture such as spatial regularity, smoothness, contrast, and randomness. We compute them across multiple orientations $\theta$ and distances $d$ to assess rotation-invariant properties. With the correlation, we can capture spatial dependencies (due to SAR image geometry acquisition), while contrast allows us to evaluate the dynamic range of SAR Rayleigh distributions, with high values for dark pixels (e.g., water) and low values for bright pixels (e.g., buildings). Moreover, the entropy captures the randomness in the texture, and the homogeneity quantifies the smoothness and uniformity of the texture.
@@ 5.1 Importance of the noise offset
As we know, SAR imagery has heavy-tailed amplitude distributions and high contrast between bright scatterers (e.g., buildings, ships) and low-reflectivity regions (e.g., calm water), we introduce a small noise offset during the forward diffusion process to add more stochasticity and help the model better capture the full dynamic range.
At each training step, the standard Gaussian noise $\varepsilon\sim\mathcal{N}(0,1)$ is perturbed by an additional random term, defined as:
@@
This modification effectively shifts the noise distribution to $\mathcal{N}(0,1+\gamma^{2})$, introducing a variability per sample and per channel without altering the spatial structure of the noise.
@@
Figure 5: Comparison of trainings - at epoch 8 - with and without noise offset and LoRA (r: rank, a: alpha).
@@
Figure 7: Comparison of image generated (1024x1024px at 40cm) during training - at epoch 8 - with the same prompt: "A satellite view of a port with a boat in the water and a forest nearby." (see more examples in Appendix  B)
As shown in Figure 7, without the noise offset, the generated images have lower contrast and a gray pixel distribution, which makes it difficult to capture important SAR-specific features, such as the contrast between land and sea. On the other hand, applying a small noise offset improves the dynamic range and enhances the physical realism of the textures, as seen in both the generated samples and the reduction in KL divergence (Table 6).We also notice that the noise offset affects the pixel distribution learning but does not alter the overall scene composition. Based on these results, we apply a noise offset by default in all subsequent training runs.
@@ 5.2 Study on the UNet, TE1 and TE2
The Text Encoders (TEs) and the UNet backbone have different roles, and understanding their contributions is essential for effective fine-tuning. To preserve the pre-existing language knowledge that is not specific to SAR, we primarily focus on the Text Encoders (TEs). Indeed, to generate complex environmental SAR scenes, such as urban, forest, and coastal areas, we use the TEs’ understanding of language, spatial relationships, and object interactions.
In Table 2, we compare different fine-tuning strategies, for the UNet backbone, Text Encoder 1 (TE1), and Text Encoder 2 (TE2), to assess their relative importance and degree of independence. Each module is either fully fine-tuned (i.e., all weights are updated) or fine-tuned using Low-Rank Adaptation (LoRA) adapters.
By default, LoRA adapters are applied to the attention projection layers of the UNet: $q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},\text{out}_{\text{proj}}]$. For the text encoders, LoRA is also applied to the same types of projection layers. Current libraries do not support LoRA on normalization layers; however, for the configuration identified as super-bowl-2, we include additional convolutional LoRA adapters in layers $\text{to}_{q},\text{to}_{k},\text{to}_{v},\text{to}_{\text{out}}.0]$ as well as convolutional layers $\text{conv}_{1},\text{conv}_{2}]$ to analyze the contribution of convolutions to generative performance.
@@
Refer to caption Refer to caption Real SAR Image Real image close up view Refer to caption Refer to caption soleil-up-7 soleil-up-7 close up view Refer to caption Refer to caption smile-road-5 smile-road-5 close up view
Figure 8: Comparison of real and generated images (1024x1024px at 40cm) - at epoch 8 - with the same prompt: "A satellite view of a dynamic city with several buildings and a network of roads."
As shown in Table 2, full fine-tuning of the UNet consistently improves alignment and realism, as reflected in lower CLIP rank mean and variance. Although the model smile-road-5 achieves the best KL divergence score—indicating good statistical similarity to real SAR distributions—its high rank variance and weak alignment suggest instability. Indeed, visual inspection (Figure 8) reveals that this model sometimes generates unrealistic features, such as handwritten-like artifacts, which do not match the prompt. This also impacts the general understanding of scene composition of the city compared to soleil-up-7.
This instability may be attributed to the fact that Text Encoder 2 (TE2) was frozen during training. As mentioned in Table  1, TE2 provides a global caption embedding that complements the token-level embeddings from Text Encoder 1 (TE1). Since TE2 was not trained, the model may have lacked the full integration of the global context provided by the text embeddings from TE2, leading to the observed inconsistencies and unrealistic features in the generated images.
As shown in Appendix B, Figure B.1, when we fully fine-tune all components of the model (mummy-pen-8), we observe that the model does not converge, and color artifacts persist even after 8 epochs. In contrast, we see a large difference between all configurations when we fully train the UNet with a LoRA on both Text Encoders (soleil-up-7), compared to the other configurations. This impacts both scene composition (with more details in the forest, realistic buildings in the city, and patterns of real mountains), resulting in more coherent and accurate scene generation. Quantitative results in Table 2 show also a lower CLIP rank and good KL distance.
To better understand which parts of the UNet contribute most to learning, we analyze the magnitude of parameter updates relative to the pretrained weights. As shown in Appendix A (Figure A.1), the largest changes are observed in the first ResNet blocks of the downsampling path and the final layers of the upsampling path. This suggests that early convolutional layers are critical for encoding SAR-specific noise and structure, further justifying the need for full fine-tuning of the UNet.
5.3 Effect of LoRA Rank and Scaling
We study how the LoRA rank $r$ and scaling factor $\alpha$ affect adaptation. With LoRA-based approach, the frozen base weight $W\in\mathbb{R}^{m\times n}$ of the generative foundation model is updated additively:
$W^{\prime}\;=\;W+\Delta W,\qquad\Delta W\;=\;\frac{\alpha}{r}\,AB$ (9)
with trainable low-rank factors $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$. By construction $\mathrm{rank}(\Delta W)\leq r$, so $r$ controls the update capacity (degrees of freedom) and the adapter parameter count $r(m+n)$, while the ratio $\alpha/r$ sets the effective update strength. In practice, $A$ is randomly initialized and $B$ is often initialized to zero so that $AB=0$ at the start of training.
We set $\alpha=r/2$, keeping $\alpha/r=0.5$ fixed. This allows us to vary $r$ to study capacity without changing the overall update magnitude. Intuitively, small $r$ may underfit (too few degrees of freedom), whereas very large $r$ may overfit and erase pretrained priors.
In our experiments, we fully fine-tune the UNet (F) and apply LoRA to both text encoders (TE1/TE2) with different $(r,\alpha)$ pairs under this rule. Results are summarized in Table 3.